# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

### v3

@@ 7 Results
Through the experiments, we aim to answer the following research questions: (1) How does PDDLLM perform relative to other LLM-based planners? (2) Can PDDLLM generalize to unseen, more complex tasks? (3) Does PDDLLM derive high-quality domains with performance comparable to expert designs? (4) How does PDDLLM’s token cost compare to other LLM-based planners? (5) How stable is PDDLLM when faced with suboptimal demonstration?
@@
We further compare PDDLLM’s planning ability with more powerful reasoning LLMs (OpenAI-o1 and DeepSeek-R1) in Table˜2. While reasoning-based models exhibit strong planning capabilities, their high computational cost prevents them from completing within the 50-second window. To sufficiently compare the planning capabilities, we extend the time limit to 500 seconds and evaluate them on the most challenging tasks. Although o1-TAMP and R1-TAMP show remarkable improvement compared to GPT-4o-based LLMTAMP, they remain less robust in long-horizon planning and fail to match PDDLLM’s performance. In contrast, our method, relying solely on GPT-4o, consistently achieves higher success rates across complex tasks. Detailed time costs for reasoning models are provided in Section˜A.6.
@@
The experimental results highlight PDDLLM’s ability to handle increasing complexity in both planning and domain derivation, consistently outperforming baseline methods. In terms of planning complexity, as shown in Figure˜3(left), PDDLLM maintains robust planning performance as task complexity grows, achieving high success rates even in scenarios involving up to 20 objects. As shown in Table˜1, performance degradation is observed in more challenging tasks such as rearrangement, where longer action sequences are required and there are more complex motion constraints. From the perspective of domain derivation complexity, PDDLLM remains effective even in tasks demanding the generation of over 100 predicates. However, success rates drop in the most complex domains, such as bridge building, primarily due to missing supporting predicates. More details of the generalization of each task are provided in section˜A.5.
@@
We evaluate the quality of planning domains generated by PDDLLM by comparing the percentage of missing or redundant predicates(actions) and the planning success rate against expert-designed domains. The full definition of missing or redundant can be found in Section˜A.11. None of the generated domains are missing actions. The predicate statistics are reported in Table˜4. The table summarizes the percentage of missing or redundant predicates, omitting tasks without such issues. These results show that PDDLLM produces high-quality domains with minimal errors, even in complex scenarios. While a few predicates may be absent, the overall logical structure remains sound, as reflected in the high overall planning success rate of 93.3% shown in Table˜1, closely matching the performance of expert-crafted domains.
@@
We compare the token efficiency of our method against LLM-based baselines on the three most complex tasks, as shown in Table˜2. Although the GPT-4o-based LLMTAMP incurs lower total token costs, it performs poorly across all three tasks. Compared to o1-TAMP and R1-TAMP, PDDLLM uses significantly fewer tokens while consistently achieving better performance. These results also underscore the challenge of deploying reasoning LLMs on real robot systems, given their high token consumption. Beyond the main research questions, we also evaluated the robustness of our method under variations in prompting styles. Results show that PDDLLM remains stable across different prompts. Full details are provided in Section˜A.7.
@@
Q5. Suboptimal demonstrations:
Lastly, we conducted experiments on the bridge-building task to evaluate how PDDLLM adapts to sub-optimal demonstrations by intentionally masking or perturbing input trajectories. The full skill set required for this task consists of pick, stack, and align actions. In the first experiment, only a pick trajectory was provided; subsequent demonstrations incrementally added stack and then align. As shown in Table˜4, success rates improved as demonstrations covered more required actions, reflecting the system’s ability to aggregate knowledge across trajectories. In a fourth experiment, we introduced a noisy trajectory corresponding to an unrelated unstack action. Although this action was initially incorporated into the domain, its negative impact was light: it was never invoked in a solution and was later pruned due to low use frequency.
@@
Stack 0.0% 12.5%
Burger Cooking 22.2% 0.0%
Bridge Building 22.2% 0.0%
@@ 8 Real Robot Deployment
We validate PDDLLM on real robots to demonstrate cross-platform deployability. The system is tested on the Agilex Piper, Franka Panda, and the UR5e, using ArUco markers for pose estimation and ROS2 for control. All platforms successfully complete tasks, including stacking, bridge building, burger cooking, and the Tower of Hanoi through direct deployment of PDDLLM (details in Section˜A.1). We also tested how PDDLLM integrates with learned skills such as VLA: logical actions serve directly as prompts to condition policies, bypassing explicit motion constraints. A real-robot experiment with VLA demonstrates this seamless interface (details in Section˜A.1.3), highlighting the framework’s flexibility across classical and learning-based control.
@@ 9 Limitation
A limitation of PDDLLM is its occasional omission of complex predicates, leading to lower success rates in harder domains like Burger Cooking and Bridge Building (Table˜1). For example, in bridge building the surface may be assembled before the base is complete, making subsequent placements infeasible. Expert-designed domains address this by introducing a predicate (all_base_finished) to enforce ordering. Recent work has sought to overcome this limitation by performing domain refining (Liang et al., 2024; Huang et al., 2025). As with many TAMP methods, modeling complex geometric conditions such as deformable objects is challenging in logical space. Nevertheless, we demonstrate that PDDLLM can be integrated with approaches targeting such problems, including VLA methods.
@@ 10 Conclusion
This paper presents PDDLLM, the first approach in the field to generate a complete planning domain from scratch, without relying on any predefined predicates or actions. By extracting logical structures directly from pre-trained LLMs, PDDLLM autonomously derives both predicates and actions, enabling fully automated domain construction. Evaluated across a wide range of environments, PDDLLM demonstrates high quality in domain derivation and strong generalizability across diverse task categories. Moreover, when integrated with the LoCA framework, PDDLLM fully automates the integration between the PDDL planning domain and the low-level motion planner. This level of automation significantly improves usability and positions the framework as an adaptable and scalable solution for robotic planning and decision-making. Compared to existing methods, PDDLLM outperforms other LLM-based baselines and closely matches the performance of expert-designed planning domains, particularly in complex and long-horizon planning scenarios. Future work will focus on integrating perception into the system, enabling domain derivation from raw sensory inputs. From the aspect of broader impact, PDDLLM lowers the barrier to deploying robots by enabling domain derivation from demonstrations, making automation more accessible and adaptable.
Appendix A Appendix
A.1 Real Robot Experiment
A.1.1 Franka Panda Arm
@@ A.1.3 VLA Experiment with UR5e
A.2 Baseline Implementation Details
LLMTAMP: LLM-based task and motion planning (LLMTAMP) builds on methods from (Huang et al., 2022; Li et al., 2022), which use pre-trained LLMs for task planning. We formulate the planning problem as a natural language description. Human demonstrations are interpreted as action sequences required to accomplish the task. The LLM then generates high-level actions to achieve the goal, which are refined into motion plans using predefined skills.
LLMTAMP-FF: LLMTAMP-FF, following the method by (Huang et al., 2023; Chen et al., 2023), extends LLMTAMP with a failure feedback loop. Upon execution failure, the system feeds the failure signal to the LLM to regenerate the task plan, repeating until success or the time limit is reached.
LLMTAMP-FR: Following Wang et al. (2024b), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM. We design a reasoner that generates detailed explanations for plan failures and incorporates them into the prompt as feedback. The LLM performs failure reasoning and then regenerates the plan accordingly.
@@ A.2 Baseline Implementation Details
A.3 Prompt template
A.3.1 Predicate imagination
@@ A.3.3 LLMTAMP baseline
A.3.4 LLMTAMP+Failure Feedback baseline
@@
A.3.5 LLMTAMP+Failure Reasoning baseline
@@ A.3.6 RuleAsMem
A.4 Experiment Tasks
A.4.1 Stacking
@@ A.4.9 Burger cooking
A.5 Generalization Across Task Complexity
@@
A.6 Time Cost of LLMTAMP Reasoning Model Variants
@@
As shown in Table˜5, our method consistently yields lower average planning times across all tasks compared to the LLMTAMP reasoning model variants. This improvement is primarily attributed to PDDLLM’s ability to structurally summarize the reasoning process into a standardized planning domain during a one-time offline inference step. As a result, no additional reasoning is required at test time. In contrast, LLMTAMP variants conduct reasoning independently for each task instance, leading to significantly higher computational costs. Such overhead makes these models impractical for deployment on physical robots, where real-time planning capabilities are often essential. As shown in Table˜2, while o1-TAMP achieves comparable performance in certain tasks, the substantially higher time cost undermines its practical value, particularly in real robotic scenarios.
@@
A.7 Prompt Variation Test
@@
A.7.1 Parallel prompting for predicate imagination
In this prompt variation, we perform parallel prompting of the LLM multiple times to obtain diverse responses simultaneously. These outputs are subsequently analyzed and aggregated to synthesize an optimal solution. While the core prompting procedure follows the same structure as described in Section˜A.3.1, additional post-processing steps are applied to summarize and consolidate the results. For prompt naming, we further prompt the LLM to select the most suitable name from among the parallel outputs. The prompt template is shown as the following:
@@ A.7.1 Parallel prompting for predicate imagination
A.7.2 Altering the prompting sequence of simulation outcomes
In this prompt variation, we alter the ordering of dimensions presented to the LLM. While the prompt template remains similar to that described in Section˜A.3.1, we do not adhere to a fixed dimension sequence. Instead, we shuffle the order of dimensions within the prompt to evaluate whether this affects the generated output. Some examples are provided here.
A.7.3 Template tuning
@@
The results, presented in Table˜6, demonstrate that our method remains robust across different prompting styles. Regardless of the prompt variation, the generated planning domains consistently solve the test tasks with a success rate approaching 100%. Minor fluctuations are attributed to randomness in the planning search process and occasional motion execution failures.
@@
A.8 Planning Time Limit Variation
@@
As outlined in the experimental design section, we set a default planning time limit of 50 seconds to reflect the real-time constraints commonly imposed on robotic systems. Although prior studies typically allow planning times on the order of minutes (Silver et al., 2020; Khodeir et al., 2023; Huang et al., 2025), the exact limits vary considerably. To assess the robustness of our approach under different time allowances, we evaluate performance at time limits of 25, 50, and 100 seconds. The results show that our method consistently outperforms all LLM-based baselines across all tested time settings and task types. Moreover, the planning domains derived by PDDLLM yield performance comparable to that of expert-designed domains under all time settings and across all tasks. All simulations were conducted on a system with an Intel Core i9-14900KF CPU without GPU acceleration, and all LLM prompting was performed using API services.
A.9 Example of PDDLLM imagined predicates
@@
Table 10: Extended examples of imagined predicates for real-world domains.
Predicate Name
@@
Collect apple of different size (small-apple ?a): 100 < w < 166.7 g
@@
Fulfill a 15-cm coffee cup (needs-refill ?a): 0 < level < 3.75 cm
(half-full ?a): 3.75 < level < 7.5 cm
@@
Measuring high school children height (medium-height ?a): 150 < h < 160 cm
@@
A.10 Test Problem Set Design
A.10.1 Dimensions of Difficulty
Task Diversity: To ensure robustness and broad applicability, we sampled over 1,200 planning tasks across nine distinct environments, including stacking, unstacking, rearrangement, alignment, parts assembly, color classification, burger cooking, and bridge building. The details of the tasks are provided in Section˜A.4. Each task category includes tasks of 3 to 20 objects, with 10 distinct tasks sampled for each object count. The resulting task plan lengths ranged from 6 to 510 steps, reflecting a wide spectrum of planning horizons. Our experiments spanned over 150 unique predicates.
Task Complexity: In task design, we consider two types of complexity: domain derivation complexity and planning complexity. Domain derivation complexity is determined by the number of predicates imagined and actions invented by PDDLLM; the more predicates and actions, the higher the complexity. Planning complexity is influenced by both the planning domain and the task. Given $n$ objects, a task plan of length $l$, and $m$ actions in the domain, the branching factor at each step is $m\times n$, resulting in a approximate complexity of $(m\times n)^{l}$. In Section˜A.10, we layout a table showcase the planning complexity (in approximate order of magnitude) of the most difficult problem in the category and the domain derivation complexity. The Tower of Hanoi task exhibits the greatest planning complexity, while bridge building presents the highest domain derivation complexity. In contrast, stack and unstack are simpler in both complexity measures.

### v4

@@ 7 Results
Through the experiments, we aim to answer the following research questions: (1) How does PDDLLM perform relative to other LLM-based planners? (2) Can PDDLLM generalize to unseen, more complex tasks? (3) Does PDDLLM derive high-quality domains with performance comparable to expert designs? (4) How does PDDLLM’s token cost compare to other LLM-based planners? (5) How does PDDLLM recover from errors such as by suboptimal demonstrations?
@@
We further compare PDDLLM’s planning ability with more powerful reasoning LLMs (OpenAI-o1 and DeepSeek-R1) in Table˜2. While reasoning-based models exhibit strong planning capabilities, their high computational cost prevents them from completing within the 50-second window. To sufficiently compare the planning capabilities, we extend the time limit to 500 seconds and evaluate them on the most challenging tasks. Although o1-TAMP and R1-TAMP show remarkable improvement compared to GPT-4o-based LLMTAMP, they remain less robust in long-horizon planning and fail to match PDDLLM’s performance. In contrast, our method, relying solely on GPT-4o, consistently achieves higher success rates across complex tasks. Detailed time costs for reasoning models are provided in Section˜B.6.
@@
The experimental results highlight PDDLLM’s ability to handle increasing complexity in both planning and domain derivation, consistently outperforming baseline methods. In terms of planning complexity, as shown in Figure˜3(left), PDDLLM maintains robust planning performance as task complexity grows, achieving high success rates even in scenarios involving up to 20 objects. As shown in Table˜1, performance degradation is observed in more challenging tasks such as rearrangement, where longer action sequences are required and there are more complex motion constraints. From the perspective of domain derivation complexity, PDDLLM remains effective even in tasks demanding the generation of over 100 predicates. However, success rates drop in the most complex domains, such as bridge building, primarily due to missing supporting predicates. More details of the generalization of each task are provided in section˜B.5.
@@
We evaluate the quality of planning domains generated by PDDLLM by comparing the percentage of missing or redundant predicates(actions) and the planning success rate against expert-designed domains. The full definition of missing or redundant can be found in Section˜B.11. None of the generated domains are missing actions. Table˜4 reports the percentage of missing and redundant predicates for the four evaluated tasks, measured against human-designed domains. The number of parallel prompts is fixed at 10, and each task is evaluated over three runs; the reported values are the averages across these runs. These results show that PDDLLM produces high-quality domains with minimal errors, even in complex scenarios. While a few predicates may be absent, the overall logical structure remains sound, as reflected in the high overall planning success rate of 93.3% shown in Table˜1, closely matching the performance of expert-crafted domains.
@@
We compare the token efficiency of our method against LLM-based baselines on the three most complex tasks, as shown in Table˜2. Although the GPT-4o-based LLMTAMP incurs lower total token costs, it performs poorly across all three tasks. Compared to o1-TAMP and R1-TAMP, PDDLLM uses significantly fewer tokens while consistently achieving better performance. These results also underscore the challenge of deploying reasoning LLMs on real robot systems for long-term use due to high token consumption. Unlike other LLM-based planners, PDDLLM spends tokens only during domain derivation; execution is handled by a PDDL solver with no additional token cost. This makes it well suited for long-term deployments that repeatedly perform similar tasks.
@@
Q5. Domain Correction:
Domain refinement can occur by adding demonstrations or refining language guidance. To evaluate how PDDLLM corrects an imprecise domain, we conducted bridge-building experiments that require coordinated picking, stacking, and alignment skills, using demonstrations that were intentionally masked or perturbed. With only a picking demo, PDDLLM had no evidence that stacking or alignment mattered, so their predicates were overlooked or collapsed into vague forms. As additional demonstrations were provided, PDDLLM progressively reconstructed the missing structure and sharpened constraints.Table 4 shows success rates rising as the missing information is supplied, demonstrating effective domain repair. We also injected a noisy, unrelated unstacking demo; it briefly appeared in the domain but was later pruned due to low usage frequency.
In addition to supplying more demonstrations, providing clearer task descriptions also improves predicate precision by offering richer contextual cues about the task objective. We include an experiment validating this effect in Section˜B.18.
@@
Stack 4.2% 8.3%
Burger Cooking 22.2% 3.7%
Bridge Building 22.2% 3.7%
@@ 8 Real Robot Deployment
We validate PDDLLM on real robots to demonstrate cross-platform deployability. The system is tested on the Agilex Piper, Franka Panda, and the UR5e, using ArUco markers for pose estimation and ROS2 for control. All platforms successfully complete tasks, including table-top stacking, bridge building, burger cooking, and the Tower of Hannoi through direct deployment of PDDLLM (details in Section˜B.1). We also evaluated how PDDLLM integrates with learned skills such as VLA on a tabletop stacking task with a UR5e. In this experiment, logical actions serve directly as prompts to condition policies,, bypassing explicit motion constraints. The real-robot experiment with VLA (details in Section˜B.1.3) demonstrates PDDLLM’s compatibility with both classical and learning-based control.
@@ 9 Limitation
A limitation of PDDLLM is its occasional omission of highly complex predicates, which can reduce problem-solving efficiency and lead to lower success rates in more difficult tasks such as Burger Cooking and Bridge Building (Table˜1). For example, in bridge building, the system may attempt to assemble the surface before the base is complete, making subsequent placements infeasible and requiring backtracking. Expert-designed domains address this by introducing predicate (all_base_finished) to enforce correct ordering. While the absence of such predicates does not directly cause planning failure, it slows down planning as the planner must perform additional feasibility checks and degrade performance under a fixed time budget. Recent work has attempted to mitigate this issue by refining planning domains through interactive or environment-feedback-based methods (Liang et al., 2024; Huang et al., 2025; Zhu et al., 2025). As with many TAMP systems, perception errors and the representation of complex dynamics and geometries, such as deformable objects and fluids, remain challenging for PDDLLM. Thus, we demonstrate that PDDLLM can be combined with approaches designed for such settings, including VLA-based manipulation methods.
10 Future works
There are a few future work directions we found meaningful to our approach. (1) Integrating perception to enable domain derivation from raw sensory inputs. Symbolic planning struggles to ground raw visual inputs into logical representations. Learning visual features directly would allow the system to derive predicates from sensor data, improving robustness to perception noise and reducing reliance on perception engineering. (2) Enabling direct interaction with the environment allows the system to obtain feedback that helps complete missing predicates, improve domain quality, reduce reliance on simulation, and better capture complex real-world dynamics. (3) Partial observability environments that requires active explore to acquire missing information.
@@ 11 Conclusion
This paper presents PDDLLM, the first approach in the field to generate a complete planning domain from scratch, without relying on any predefined predicates or actions. By extracting logical structures directly from pre-trained LLMs, PDDLLM autonomously derives both predicates and actions, enabling fully automated domain construction. Evaluated across a wide range of environments, PDDLLM demonstrates high quality in domain derivation and strong generalizability across diverse task categories. Moreover, when integrated with the LoCA framework, PDDLLM fully automates the integration between the PDDL planning domain and the low-level motion planner. This level of automation significantly improves usability and positions the framework as an adaptable and scalable solution for robotic planning and decision-making. Compared to existing methods, PDDLLM outperforms other LLM-based baselines and closely matches the performance of expert-designed planning domains, particularly in complex and long-horizon planning scenarios.
Reproducibility Statement
To facilitate reproduction of our results, we provide clear references to the relevant sections of the paper. The overall planning pipeline is described in Section˜4. The extraction of the PDDL domain from demonstrations is illustrated in Figure˜1, with additional details provided in Section˜B.12. The procedure for predicate imagination and action invention is discussed in the Section˜5, where we also outline the prompt templates and prompting strategies in greater detail in Section˜B.3. Finally, the LoCA framework is described comprehensively in Section˜5.2, and experiment details shown in Section˜8 and Section˜B.1.
Ethics Statement
We affirm our commitment to the principles laid out in the ICLR Code of Ethics, including honesty, fairness, transparency, and avoidance of harm. The human demonstrations collected do not contain identifying information. Our research does not use sensitive personal or private data. We have carefully documented our methodology, data sources, and limitations to promote reproducibility and accountability. We see no foreseeable misuses of the methods or results discussed in this work, but we remain alert to potential downstream impacts and encourage continued scrutiny in the broader research community.
Appendix A The Use of Large Language Models (LLMs)
This paper evaluates LLMs within the robotics domain, where LLMs serve as the testing target of our method. Beyond this role, LLMs were used as grammar-checking and text-polishing tools to improve readability (e.g., grammar, clarity, and style). They were not involved in the ideation, experimental design, analysis, or interpretation of results, and did not contribute to the scientific content. The authors take full responsibility for the entirety of the work presented.
Appendix B Appendix
B.1 Real Robot Experiment
Ten runs of each real robot experiment shown in Section 8 are performed. The success rates of real-world deployment on different robot platforms are shown as below:
Table 5: Real-world success rate across tasks.
Metric Tower of Hanoi (Franka) Bridge Building (Franka) Burger (Piper) Table-top Stacking (UR5e)
Success Rate 9/10 8/10 7/10 7/10
B.1.1 Franka Panda Arm
@@ B.1.3 VLA Experiment with UR5e
B.2 Baseline Implementation Details
LLMTAMP: LLM-based task and motion planning (LLMTAMP) builds on methods from (Huang et al., 2022a; Li et al., 2022), which use pre-trained LLMs for task planning. We formulate the planning problem as a natural language description. Human demonstrations are interpreted as action sequences required to accomplish the task. The LLM then generates high-level actions to achieve the goal, which are refined into motion plans using predefined skills.
LLMTAMP-FF: LLMTAMP-FF, following the method by (Huang et al., 2022b; Chen et al., 2024a), extends LLMTAMP with a failure feedback loop. Upon execution failure, the system feeds the failure signal to the LLM to regenerate the task plan, repeating until success or the time limit is reached.
LLMTAMP-FR: Following Wang et al. (2024), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM. We design a reasoner that generates detailed explanations for plan failures and incorporates them into the prompt as feedback. The LLM performs failure reasoning and then regenerates the plan accordingly.
@@ B.2 Baseline Implementation Details
B.3 Prompt template
B.3.1 Predicate imagination
@@ B.3.3 LLMTAMP baseline
B.3.4 LLMTAMP+Failure Feedback baseline
@@
B.3.5 LLMTAMP+Failure Reasoning baseline
@@ B.3.6 RuleAsMem
B.4 Experiment Tasks
B.4.1 Stacking
@@ B.4.9 Burger cooking
B.5 Generalization Across Task Complexity
@@
B.6 Time Cost of LLMTAMP Reasoning Model Variants
@@
As shown in Table˜6, our method consistently yields lower average planning times across all tasks compared to the LLMTAMP reasoning model variants. This improvement is primarily attributed to PDDLLM’s ability to structurally summarize the reasoning process into a standardized planning domain during a one-time offline inference step. As a result, no additional reasoning is required at test time. In contrast, LLMTAMP variants conduct reasoning independently for each task instance, leading to significantly higher computational costs. Such overhead makes these models impractical for deployment on physical robots, where real-time planning capabilities are often essential. As shown in Table˜2, while o1-TAMP achieves comparable performance in certain tasks, the substantially higher time cost undermines its practical value, particularly in real robotic scenarios.
@@
B.7 Prompt Variation Test
@@
B.7.1 Parallel prompting for PDDLLM
We perform parallel prompting of the LLM multiple times to obtain multiple responses simultaneously. These outputs are subsequently analyzed and aggregated to synthesize an optimal solution. The core prompting procedure follows the same structure as described in Section˜B.3.1. All the domains failed to execute are removed from the candidate pool. Among those successful planning domain candidates, additional post-processing steps are applied to select, summarize and consolidate the results. For predicate naming, we further prompt the LLM to select the most suitable name from among the parallel outputs. For predicate selection, the prompt template is shown as the following:
@@ B.7.1 Parallel prompting for PDDLLM
Similarly, for optimal actions selection:
Here is a list of actions describing the same robot skill: action list. Please choose the PDDL action from the provided ones to best describe the scenario. Return the chosen one in PDDL syntax.
@@
B.7.2 Altering the prompting sequence of simulation outcomes
In this prompt variation, we alter the ordering of dimensions presented to the LLM. While the prompt template remains similar to that described in Section˜B.3.1, we do not adhere to a fixed dimension sequence. Instead, we shuffle the order of dimensions within the prompt to evaluate whether this affects the generated output. Some examples are provided here.
B.7.3 Template tuning
@@
The results, presented in Table˜7, demonstrate that our method remains robust across different prompting styles. Regardless of the prompt variation, the generated planning domains consistently solve the test tasks with a success rate approaching 100%. Minor fluctuations are attributed to randomness in the planning search process and occasional motion execution failures.
@@
B.8 Planning Time Limit Variation
@@
As outlined in the experimental design section, we set a default planning time limit of 50 seconds to reflect the real-time constraints commonly imposed on robotic systems. Although prior studies typically allow planning times on the order of minutes (Silver et al., 2021; Khodeir et al., 2023; Huang et al., 2025), the exact limits vary considerably. To assess the robustness of our approach under different time allowances, we evaluate performance at time limits of 25, 50, and 100 seconds. The results show that our method consistently outperforms all LLM-based baselines across all tested time settings and task types. Moreover, the planning domains derived by PDDLLM yield performance comparable to that of expert-designed domains under all time settings and across all tasks. All simulations were conducted on a system with an Intel Core i9-14900KF CPU without GPU acceleration, and all LLM prompting was performed using API services.
B.9 Example of PDDLLM imagined predicates
@@
Table 11: Examples of predicate generation across different task and scenes.
@@
Collect apples of different sizes (small-apple ?a): 100 < w < 166.7 g
@@
Fulfill a 15-cm coffee cup (needs-refill ?a): 0 < level < 7.5 cm
@@
Measure high school students’ heights (medium-height ?a): 150 < h < 160 cm
@@
Predicates for Temperature
Heat a kitchen stove (stove-hot ?a): 110 < T < 200 °C
Boil milk (cold-milk ?a): 0 < T < 10 °C
(warm-milk ?a): 20 < T < 40 °C
(boiling-milk ?a): 90 < T < 100 °C
B.10 Test Problem Set Design
B.10.1 Dimensions of Difficulty
Task Diversity: To ensure robustness and broad applicability, we sampled over 1,200 planning tasks across nine distinct environments, including stacking, unstacking, rearrangement, alignment, parts assembly, color classification, burger cooking, and bridge building. The details of the tasks are provided in Section˜B.4. Each task category includes tasks of 3 to 20 objects, with 10 distinct tasks sampled for each object count. The resulting task plan lengths ranged from 6 to 510 steps, reflecting a wide spectrum of planning horizons. Our experiments spanned over 150 unique predicates.
Task Complexity: In task design, we consider two types of complexity: domain derivation complexity and planning complexity. Domain derivation complexity is determined by the number of predicates imagined and actions invented by PDDLLM; the more predicates and actions, the higher the complexity. Planning complexity is influenced by both the planning domain and the task. Given $n$ objects, a task plan of length $l$, and $m$ actions in the domain, the branching factor at each step is $m\times n$, resulting in a approximate complexity of $(m\times n)^{l}$. In Section˜B.10, we layout a table showcase the planning complexity (in approximate order of magnitude) of the most difficult problem in the category and the domain derivation complexity. The Tower of Hanoi task exhibits the greatest planning complexity, while bridge building presents the highest domain derivation complexity. In contrast, stack and unstack are simpler in both complexity measures.